Use the table below to determine the MPC and MPS.

Disposable Income

Consumption

Saving

$1000

$1100

-$100

2000

1600

   400

3000

2100

   900

1. Using the above information, what is the MPC and MPS when the DI is 3000?

MPC =                                                MPS =

2. What equation could you use to determine the Multiplier, using MPC and MPS?

Multiplier =

Multiplier =

3. If there is an initial investment spending of $5,000; what would the total change in GDP be?
Answer #1

1. MPC = Change in Consumption / Change in Disposable income

Change in Consumption = 2100 - 1600 = 500

Change in income = 3000 - 2000 = 1000

MPC = 500/1000 = 0.5

MPS = 1 - MPC = 1 - 0.5 = 0.5

2. Multiplier = 1/(1 - MPC) = 1/(1 - 0.5) = 1/0.5 = 2

Multiplier = 1/MPS = 1/0.5 = 2

3. Change in GDP = Change in Investment spending / (1 - MPC)

Change in GDP = 5000/0.5 = 10,000